How do we calculate the material performance index of a low-cost column? The height does not matter but the cross-sectional diameter must be 10 +/- 0.1 cm. (Cross section is circular, column must support a 1-ton axial force without buckling)
Answer #1

Solution :

P = 0.4 fckAc+0.67fyAsc

P = 0.4fck (Ag-Asc) + 0.67fyAsc

Ag = pi/4 x 1002 = 7.85 x 103 mm2

(10x103) = 0.4 x20 (7.85x103 - Asc) + 0.67 x 415 x Asc

= 195.51 mm2Critical buckling load: El (KL)2 Effective length factor K Le 2L Le L L0.5L PIN-FIXE FNDS IN PIN FNDら XED FREE IXED-FIXED ENCritical Buckling Stress of a Column Formula 2 Tt E 2 σ-critical or allowable stress of column in Pa E-young modulus (elastic

Buckling can be defined as the sudden large deformation of structure due to a slight increase of an existing load under which the structure had exhibited little, if any, deformation before the load was increased.”

For M25

= 246740.11/900 = 274.15